OpenStudy (theraggedydoctor):

If \(b^{2x+1}=b^{3x-1}\) for all values of \(b\), what is the value of \(x\)?

OpenStudy (theraggedydoctor):

The first one is 2x+1 and the second one is 3x-1.

OpenStudy (amistre64):

since the bases are the same, then they must be equal when the exponents are equal

OpenStudy (theraggedydoctor):

What?

OpenStudy (amistre64):

its there for you to read, im not going to resay it ....

OpenStudy (theraggedydoctor):

I already read it..

OpenStudy (amistre64):

what?

OpenStudy (theraggedydoctor):

I dont get what you meant.

OpenStudy (amistre64):

what part of it is a mystery? bases, exponents, or equal to?

OpenStudy (theraggedydoctor):

Exponents.

OpenStudy (amistre64):

the ^(...) is called an exponent, or a power

OpenStudy (amistre64):

b^n = b^m only if n=m

OpenStudy (theraggedydoctor):

Oh.. okay.

OpenStudy (theraggedydoctor):

2

OpenStudy (amistre64):

correct